OpenStudy (anonymous):

factor 4x^2-25y^2z^2

OpenStudy (anonymous):

(2x+5yz)(2x-5yz) ? ? ? Is that it?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

(2x+5yz)(2x-5yz) is correct

OpenStudy (anonymous):

4x^2 - (25y^2)(z^2) note we have 2 squares. the first is 4x^2 and square root of this is 2x the 2nd is (25y^2)(z^2) made from 2 different squares. square root of 25y^2 is 5y and square root of z^2 is z so combined square root of (25y^2)(z^2) is 5yz Now you have to remember the law of the difference of two squares. which says (a^2)-(b^2)= (a-b)(a+b) if a=2x, and b=5yz then yes your answer comes out correct. (2x+5yz)(2x-5yz)